Hello, I need help solving this word problem. I don't know where to begin.
Christine went shopping and bought each of her five nephews a gift, either a video costing 14.95 or a cd costing 16.88. She spend 76.68 on gifts. How many videos and how many cds did she buy.

Answer by checkley77(12844)

14.95x+16.88(5-x)=76.68
14.95x+84.40-16.88=76.68
14.95x-16.88=76.88-84.40
-1.93x=-7.72
x=-7.52/-1.93
x=4 is the number of videos bought.
5-4=1 is the number of cds bought.
Proof:
14.95*4+16.88*1=76.88
59.80+16.88=76.68
76.68=76.68